HIGH RESOLUTION TDL SPECTROSCOPY OF THE Ar–CH_{4} COMPLEX

\underline{M. WANGLER}, D. A. ROTH, I. PAK, and \underline{G. WINNEWISSER}, I. Physikalisches Institut, Universität zu Köln, Zülpicher Str. 77, 50937 Köln, Germany; M. GELEIJNS, P.E.S. WORMER, and A. VAN DER AVOIRD, Institute of Theoretical Chemistry, University of Nijmegen, Toernooiveld 1, 6525 ED Nijmegen, The Netherlands.

The spectrum of the weakly bound complex Ar–CH_{4} in the 7 μm region was discovered, analysed, and compared with a spectrum, predicted from ab initio calculations. The measurements were made by probing a supersonic gas expansion with a tunable diode laser (TDL). Several bands of Ar–CH_{4} associated with different ro-vibrational transitions of the ν_{4} vibration of CH_{4} were recorded and analyzed in a spectral region from 1295 to 1330 cm^{−1}. In particular the following transitions were studied: j = 1 ← 0 (@ 1311 cm^{−1}) reported in [Z. Naturforsch. 53, 725 (1998)], j = 0 ← 1 (@ 1301 cm^{−1}), j = 1 ← 1 (@ 1306 cm^{−1}), j = 2 ← 1 (@ 1316 cm^{−1}), and j = 3 ← 2 transitions (@ 1322 cm^{−1}). Here, j denotes the angular momentum of the methane unit inside the complex. The experimental results are compared with ab initio calculations. The close agreement between observed and ab initio spectra is convincingly demonstrated with respect to the gross spectral features, including many details of the spectra.